### v1

# MonoDETRNext: Next-generation Accurate and Efficient Monocular 3D Object Detection Method

Abstract
[…]
1 Introduction
3D object detection boasts a vast range of applications, from autonomous driving and robotic navigation to smart surveillance and virtual reality. These applications hinge critically on the precise recognition and localization of objects within 3D space. However, many 3D object detection methods often rely on expensive equipment such as LiDAR[1, 2, 3, 4, 5, 6], significantly limiting their practical deployment. While several purely visual strategies have achieved notable success, these generally involve multiple cameras[7, 8, 9] and are still limited under certain circumstances. Furthermore, they typically require substantial computational power to operate effectively.
Monocular 3D detection methods have not been given much attention, largely due to their relatively low accuracy. In many scenarios that require 3D detection, it is feasible to employ LiDAR or multiple cameras, supported by top-tier computing resources. However, in extreme cases where a LiDAR or one of the cameras fails, these high-precision algorithms might be rendered ineffective. Similarly, in scenarios with restricted device availability or budget constraints, multi-view and LiDAR-dependent algorithms may not be viable. Therefore, researching monocular 3D detection algorithms holds practical significance.
In the realm of 2D detection, models from the DETR series[10, 11, 12, 13] currently stand unmatched. MonoDETR[14] represents the first end-to-end monocular 3D object detection model within this category. Its performance surpasses all previous monocular 3D detection approaches. However, the impressive results of MonoDETR are not sufficient to satisfy our aspirations. Introducing some of the recent advancements in technology and strategies into monocular 3D vision detection, this paper presents a new-generation model, MonoDETRNext-F.
We also observed that in MonoDETR, the acquisition of depth information is overly simplistic, which is crucial for monocular 3D detection models since depth information is one of the few available 3D data types. The accuracy of depth information is closely linked to the precision of detection. To improve depth information acquisition, we referred to established 3D depth estimation networks and designed a depth predictor that directly extracts depth information from images. We integrated this predictor with MonoDETRNext-F to develop MonoDETRNext-A, which offers enhanced detection capabilities.
Through this work, we aim to provide a significant starting point for the future development of monocular 3D vision detection models, laying a solid foundation for subsequent research and applications.
The salient innovations encapsulated within this manuscript may be delineated as follows:
  • 1)
    Proposing two novel monocular 3D object detection models, namely MonoDETRNext-F and MonoDETRNext-A, the former adeptly balancing speed and precision, while the latter accentuates precision-centric objectives.
  • 2)
    Architecting a hybrid vision encoder for 3D object detection, proficient in the efficient extraction of features crucial for monocular 3D object detection tasks, alongside the integration of a refined, yet lightweight depth estimation module, envisaged to bolster detection accuracy.
  • 3)
    Drawing upon methodologies hitherto entrenched in 2D object detection paradigms, we advance a cogent strategy for object query generation tailored to the exigencies of monocular 3D object detection, underpinned by meticulously designed training strategies, aimed at optimizing model performance efficaciously.
2 Related Work
Current 3D object detection methods can generally be classified into two categories: camera-based methods and fusion methods integrating LiDAR and other sensors.
Camera-based methods can be further divided into monocular (single-view) and multi-view methods based on the number of input viewpoints. Monocular detectors utilize only forward-facing images as input, addressing more complex tasks with limited 2D information. Multi-view detectors simultaneously encode images of the surrounding scene, leveraging relationships between viewpoints to understand 3D space. On the other hand, fusion methods based on LiDAR and other sensor integration rely on inputs from devices such as depth cameras and LiDAR, which provide a fusion of various sensor data types, including images and point clouds. Consequently, they can acquire richer and more comprehensive depth information. Despite their higher cost, these methods typically exhibit greater robustness and accuracy in complex environments, as they can exploit the advantages of multiple sensors and integrate information from different data sources.
Figure 1: Comparison of mainstream 3D detection models, with different colors representing distinct functional modules.
2.1 MonoDETR and other monocular 3D detection
MonoDETR [14] is a state-of-the-art method that leverages rendered transmittance to predict depth maps from a single RGB image. By capturing subtle cues in the input image, MonoDETR achieves improved accuracy and robustness to varying lighting conditions compared to traditional monocular depth estimation methods.
Several other monocular 3D reconstruction method have been proposed in recent years. For example, MonoDTR[15] is a deep learning model that predicts depth maps from single RGB images using a transformer-based architecture. While MonoDTR achieves high accuracy, it requires additional LiDAR data for training assistance. Meanwhile, CaDDN[16] and Monorun[17] not only require LiDAR data during training but also during inference. Autoshape[18] integrates CAD data into the model to augment the restricted 3D cues. MonoDETR requires minimal 2D-3D geometric priors and does not necessitate additional annotations. Our MonoDETRNext inherits this characteristic.
Alternative methodologies, as demonstrated by MonoDLE [19], PGD [20], and PackNet [21], integrate multi-scale feature fusion and attention mechanisms for depth map estimation and error analysis, yielding improved performance. Despite their high accuracy, these approaches incur substantial computational costs and demand significant memory resources. Conversely, MonoDETR is characterized by its lightweight and efficient nature. Furthermore, MonoDETRNext-F surpasses it in speed and efficacy, while MonoDETRNext-A demonstrates markedly superior performance.
2.2 Multi-view 3D object detection
To jointly extract features from surrounding views, DETR3D [22] initially employs a set of 3D object queries, which are then back-projected onto multi-view images to aggregate features. The PETR series [23, 24, 25] further introduces the generation of 3D positional features, avoiding unstable projections, and explores the advantages of temporal information from the preceding frame.
Alternatively, BEVFormer [7] and its improvements [8, 9] generate BEV (Bird’s Eye View) features using learnable BEV queries and introduce a spatiotemporal BEV transformer for visual feature aggregation. Subsequent research has also investigated cross-modal distillation [26, 27] and masked image modeling [28, 29] to enhance performance.
2.3 LiDAR and multi-source information fusion 3D object detection
Methods such as DeepFusion [30] and PointPainting [31] represent notable advancements in the integration of LiDAR point cloud data with camera imagery to facilitate precise object detection within three-dimensional spatial environments. This fusion strategy optimally exploits the synergies inherent in disparate sensor modalities, amalgamating spatial depth cues with color texture information, thereby fortifying the resilience and accuracy of detection outcomes.
The integration of principles from BEVFormer into fusion paradigms, exemplified by BevFusion [6], has spurred further refinements culminating in enhanced precision, as evidenced by exemplar models such as those delineated in MV2D [4] and Futr3d [32]. Recent endeavors, typified by mmFusion [33], have expanded the purview of fusion methodologies by integrating data from multiple sensors, including cameras, LiDAR, and radar, resulting in notable performance strides.
Concurrently, the domain has witnessed the emergence of large-scale architectures, exemplified by OMNI3D [34] and GLEE [35], which have showcased remarkable efficacy in 3D object detection tasks. Leveraging copious training data and sophisticated model architectures characterized by billions of parameters or more, these frameworks have been trained utilizing advanced optimization algorithms, thereby augmenting detection performance and precision.
3 Method
3.1 Model Overview
The overall frameworks of MonoDETRNext-A and MonoDETRNext-F are illustrated in Figure 2. The main difference between these two models lies in the disparity of the depth predictors, which will be elaborated in Section 3.3. In Section 3.2, the design principles of the efficient encoder and the extraction of visual and depth features will be explicated. Lastly, in Section 3.4, the meticulous generation of object queries and the loss utilized will be expounded.
Figure 2: The schematic depiction of MonoDETRNext. The distinction between MonoDETRNext-A and MonoDETRNext-F primarily resides in their respective depth prediction mechanisms. The provided illustration delineates the intricate depth prediction scheme adopted by MonoDETRNext-A, whereas the depth predictor employed in MonoDETRNext-F remains congruent with that of MonoDETR.
3.2 Efficient Hybrid Vision Encoder
[…]
Inspired by the architecture of RT-DETR[38], we introduce an innovative Efficient Hybrid Vision Encoder tailored specifically for 3d object detection tasks. This encoder is characterized by a diminished computational footprint while preserving the efficacy of feature extraction. Illustrated in Figure 2, our proposed encoder comprises two integral elements: a singular encoder layer and a Cross-Scale Feature Integration Module (CFIM) grounded in CNN. The CFIM functions as a fusion unit, amalgamating adjacent features into novel representations, as delineated in Figure 3. This fusion process is succinctly formalized by the ensuing equations:
[…]
Compared to the hybrid encoder of RT-DETR, our hybrid encoder exhibits significant differences in the Fusion module.
3.3 Accurate Depth Predictor
The principal disparity between our proposed models, MonoDETRNext-A and MonoDETRNext-F, resides in their Depth Predictor architecture. The initial Depth Predictor in MonodDETR, by design, featured a lightweight configuration comprising solely two $3$ convolutional layers, extracting $f_{D}\in\mathbb{R}^{\frac{H}{16}\times\frac{W}{16}\times(C)}$ from features preprocessed by the backbone. The accuracy of monocular 3D object detection correlates significantly with the quality of depth information obtained. In pursuit of refined depth estimation and subsequent enhancement of 3D detection precision, we undertook a redesign of the Depth Predictor. The revamped depth predictor now directly leverages the image data itself to extract depth-related features, circumventing secondary processing of backbone-derived features, thereby alleviating the potential for misalignment learning during model training.
The architectural of this novel Depth Predictor draws inspiration from several state-of-the-art monocular depth estimation models, such as lite-mono[39] and monodepth2[40]. Its specific structural configuration, as depicted in Figure 2, initiates with the processing of images of dimensions $H\times W\times 3$ through a convolutional perturbation module, where image data undergoes downsampling via a $3\times 3$ convolutional operation. Subsequently, two additional $3\times 3$ convolutional layers, each with a stride of $1$, are deployed for local feature extraction, yielding feature maps of dimensions $\frac{H}{2}\times\frac{W}{2}\times C_{1}$. In the ensuing stage, these features are concatenated with the pooled three-channel input image, followed by another downsampling step through a $3\times 3$ convolutional layer with a stride of $2$, resulting in feature maps of dimensions $\frac{H}{4}\times\frac{W}{4}\times C_{2}$. The integration of features with the pooled input image in the downsampling layers serves to mitigate the loss of spatial information attributable to the reduction in feature dimensions, drawing inspiration from ESPNetv2[41]. Subsequent to this, RGFI and SDC are introduced to facilitate the acquisition of rich hierarchical feature representations. The ensuing downsampling layers also inherit connected features from the preceding downsampling layer. Analogously, the output feature maps undergo further processing until their dimensions align with those of MonoDETR[14], specifically $\frac{H}{16}\times\frac{W}{16}\times C_{4}$.
It is worth mentioning that MonoDETRNext-A utilizes our custom-designed Depth Predictor, while MonoDETRNext-F utilizes the original MonoDETR. After obtaining $f_{D}$, the subsequent operations for both our models, as well as MonoDETR, involve encoding through a linear classification layer and an encoder layer to obtain $D_{fg}\in\mathbb{R}^{\frac{H}{16}\times\frac{W}{16}\times(k+1)}$ along with other additional data. These data are then fed into the depth-guided decoder to parse features.
The Sequential Dilated Convolution (SDC) module is introduced to extract multi-scale local features using dilated convolutions. Similar to lite-mono, we adopt a staged approach by inserting multiple consecutive dilated convolutions with varying dilation rates to effectively aggregate multi-scale context.
Given a 2D signal $x[i]$, the output $y[i]$ of a 2D dilated convolution can be defined as:
$y[i]=\sum_{k=1}^{K}x[i+r\cdot k]w[k],$ (5)
where $w[k]$ is a filter of length $K$, and $r$ denotes the dilation rate used for convolving the input $x[i]$. In standard non-dilated convolutions, $r=1$. By employing dilated convolutions, the network can maintain a fixed size for the output feature map while achieving a larger receptive field. Considering an input feature $X$ with dimensions $H\times W\times C$, the output $\hat{X}$ of our SDC module is computed as follows:
$\hat{X}=X+\text{ Linear }_{G}\left(\text{ Linear }\left(BN\left(\text{ Conv }_{r}(X)\right)\right)\right),$ (6)
where $\text{Linear}_{G}$ denotes pointwise convolution followed by the GELU[42] activation. $BN$ represents the batch normalization layer, and $\text{Conv}_{r}(\cdot)$ is a $3\times 3$ convolutional layer.
The Regional-Global Feature Interaction (RGFI) module operates as follows: given an input feature map $X$ of dimensions $H\times W\times C$, it is linearly projected into queries $Q=XW_{q}$, keys $K=XW_{k}$, and values $V=XW_{v}$, where $W_{q}$, $W_{k}$, and $W_{v}$ are weight matrices. Cross-covariance attention[43] is employed to enhance the input $X$:
$\widetilde{X}=\operatorname{Attention}(Q,K,V)+X,$ (7)
Subsequently, non-linearity is introduced to enhance the features:
$\hat{X}=X+\text{ Linear }_{G}(\operatorname{Linear}(\operatorname{LN}(\tilde{X}))),$ (8)
where $LN$ represents a layer normalization[44] operation. Four variants of the depth encoder are designed based on different numbers of channels, SDC blocks. The complete process is illustrated in Figure 3 .
Figure 3: The fusion block within CFIM is depicted, showcasing the architectures of the Sequential Dilated Convolution (SDC) module and the Regional-Global Feature Interaction (RGFI) module proposed therein.
3.4 Effective Query Generation and Overall Loss
In DETR, object queries are a set of learnable embeddings optimized by the decoder and mapped to classification scores and bounding boxes via prediction heads. However, these object queries are challenging to interpret and optimize as they lack explicit physical meaning. Subsequent works[45, 46, 47, 11, 48] have improved the initialization of object queries and extended them to include content queries and position queries (anchors). In 3D object detection, the position of objects is no longer a simple 2D bounding box but includes depth information in a 3D bounding box. Therefore, a more accurate and interpretable method for generating 3D object queries is needed. To address this, we introduce the process of generating object queries from 2D DETR models into 3D object detection, expanding it into depth-aware object queries.
To put it simply, we have designed a target query generation method based on the characteristics of 3D objects. In MonoDETR[14], the final output of the bbox is a six-dimensional vector, $[x_{c},y_{c},x,y,w,h]$, $x_{c},y_{c}$ represent the 3D center point of the object in the image coordinates, and $x,y,w,h$ define the 2D bbox. However, MonoDETR’s object query generates only a two-dimensional vector, which is clearly not ideal. MV2D[4] provided us with inspiration, demonstrating that 2D bounding boxes can enhance the performance of 3D detection. Therefore, we incorporated the process of generating position queries from the 2D DETR model into our system, treating the queries as bounding boxes for loss computation. Specifically, the process is as follows:
$\mathcal{L}_{enc}(\hat{y},y)&=\mathcal{L}_{\text{box }}(\hat{b},b)+\mathcal{L}_{cls}(\hat{c},\hat{b},y,b)\\ &=\mathcal{L}_{\text{box }}(\hat{b},b)+\mathcal{L}_{cls}(\hat{c},c,IoU)\end{aligned}$ (9)
where $\hat{y}$ and $y$ denote prediction and ground truth, $\hat{y}={\hat{c},\hat{b}}$ and $y={c,b}$, $c$ and $b$ represent categories and bounding boxes, respectively. We introduce the IoU score. Additionally, we incorporate the IoU score into the classification branch’s objective function (similar to VFL[49] and RT-DETR[38]), achieving consistency constraints for positive sample classification and localization.
For generating 3D centers, we experiment with several methods, but regretfully find that a set of learnable embeddings yields the best results. Finally, our overall loss is similar to that of MonoDETR, except $\mathcal{L}_{enc}$:
$\mathcal{L}_{\mathrm{overall}}=\frac{1}{N_{\mathrm{gt}}}\cdot\sum_{n=1}^{N_{gt}}{\left(\mathcal{L}_{2D}+\mathcal{L}_{3D}+\mathcal{L}_{enc}\right)}+\mathcal{L}_{\mathrm{dmap}}$ (10)
where $\mathcal{L}_{d\text{map }}$ represents the focal loss[50] of the predicted categorical foreground depth map $D_{fg}$ in Section.3.3 and $N_{gt}$ denotes the number of ground-truth objects. Additionally, it is noteworthy that the proportion of the total weight attributed to the $\mathcal{L}_{d\text{map}}$ loss during training differs between MonoDETRNext-A and MonoDETRNext-F, with MonoDETRNext-A having a higher percentage.
4 Experiments
4.1 Setting
4.1.1 Dataset
MonoDETRNext is assessed on the extensively utilized KITTI[51] benchmark, which encompasses 7,481 training and 7,518 test images. Following the protocol outlined in previous studies [52, 53], we segregate 3,769 validation images from the training set. Evaluation is conducted across three difficulty levels: easy, moderate, and hard. Performance metrics are computed based on average precision ($AP$) of bounding boxes in both 3D space ($AP_{3D}$) and the bird’s-eye view ($AP_{BEV}$). These metrics are assessed at 40 recall positions to ensure comprehensive evaluation.
Note: The reason for not using commonly used datasets for other types of 3D detection methods is that the vast majority of monocular 3D detection methods have only been validated on the KITTI, and we cannot find their results on other datasets.
4.1.2 Implementation details
We employed ResNet50[54] as our backbone. While we also explored more advanced networks such as ConvNext[55], regrettably, these sophisticated architectures did not yield performance improvements. We utilized eight attention heads for all attention modules, including RGFI. MonoDETRNext-F was trained for 230 epochs on a single RTX 4090 GPU, with a batch size of 8 and a learning rate of $2\times 10^{-4}$. MonoDETRNext-A was trained for 300 epochs, with the learning rate decayed between epochs 225 and 275. We employed the AdamW[56] optimizer with weight decay of $10^{-4}$. For MonoDETRNext-F, the learning rate was decreased by a factor of 0.1 at epochs 145 and 205 and for MonoDETRNext-A, the process is between 165 and 205. The number of object queries was consistent with MonoDETR, set at 50. Inference speed was evaluated on a single RTX 3090 GPU.
4.2 Comparison with SOTA
Table 1: Monocular performance of the car category on KITTI test and val sets. We utilize bold numbers to highlight the best results, and color the second-best ones expect our methods and our gain over them in blue. The FPS values are obtained through speed testing conducted on individual models using a single NVIDIA GeForce RTX 3090 GPU. During testing, a batch size of 1 was utilized.
Method $FPS_{bz=1}$ Test, $AP_{3D}$ Test, $AP_{BEV}$ Val, $AP_{3D}$
Easy Mod. Hard Easy Mod. Hard Easy Mod. Hard
SMOK[57] 17.6 14.03 9.76 7.84 20.83 14.49 12.75 14.76 12.85 11.50
MonoPair [58] - 13.04 9.99 8.65 19.28 14.83 12.89 16.28 12.30 10.42
RTM3D [59] 23.3 13.61 10.09 8.18 - - - 19.47 16.29 15.57
PatchNet [60] - 15.68 11.12 10.17 22.97 16.86 14.97 - - -
D4LCN [61] - 16.65 11.72 9.51 22.51 16.02 12.55 - - -
DDMP-3D [62] 9.4 19.71 12.78 9.80 28.08 17.89 13.44 - - -
Kinematic3D [63] 15.6 19.07 12.72 9.17 26.69 17.52 13.10 19.76 14.10 10.47
MonoRUn [17] 20.2 19.65 12.30 10.58 27.94 17.34 15.24 20.02 14.65 12.61
CaDDN [16] - 19.17 13.41 11.46 27.94 18.91 17.19 23.57 16.31 13.84
PGD [20] 23.9 19.05 11.76 9.39 26.89 16.51 13.49 19.27 13.23 10.65
MonoDLE [19] 25.0 17.23 12.26 10.29 24.79 18.89 16.00 17.45 13.66 11.68
MonoRCNN [64] 24.7 18.36 12.65 10.03 25.48 18.11 14.10 16.61 13.19 10.65
MonoGeo [65] - 18.85 13.81 11.52 25.86 18.99 16.19 18.45 14.48 12.87
MonoFlex [66] 27.9 19.94 13.89 12.07 28.23 19.75 16.89 23.64 17.51 14.83
GUPNet [67] 28.7 20.11 14.20 11.77 - - - 22.76 16.46 13.72
MonoDTR [15] 27.0 21.99 15.39 12.73 28.59 20.38 17.14 24.52 18.57 15.51
AutoShape [18] - 22.47 14.17 11.36 30.66 20.08 15.59 20.09 14.65 12.07
MonoDETR[14] 26.5 25.00 16.47 13.58 33.60 22.11 18.60 28.84 20.61 16.38
SSD-MonoDETR[68] - 24.52 17.88 15.73 33.59 24.35 21.98 29.53 21.96 18.20
MonoDETRNext-F 29.2 27.21 21.69 20.16 34.56 28.12 26.31 30.28 24.77 20.38
MonoDETRNext-A 20.8 29.94 24.14 23.79 37.32 30.68 30.29 33.44 25.75 22.40
The table 1 presents a comparison of various monocular 3D object detection models on the KITTI dataset, focusing on the car category’s performance. The evaluation metrics $AP_{3D}$ and $AP_{BEV}$ , which are both at 40 recall positions, across different difficulty levels (Easy, Moderate, and Hard) for both the test and validation sets. Additionally, the FPS metric for batch size 1 is provided to gauge the inference speed of each model.
Among the methods listed, our methods MonoDETRNext-F and MonoDETRNext-A stands out. These variants exhibit notable advantages over other models in terms of both accuracy and efficiency.
In terms of accuracy, MonoDETRNext-F achieves impressive $AP_{3D}$ and $AP_{BEV}$ scores across all difficulty levels, surpassing all other models listed. Specifically, it outperforms MonoDETR and SSD-MonoDETR in $AP_{3D}$, showcasing its superior ability to accurately detect 3D objects. Additionally, it achieves the highest $AP_{BEV}$ scores, indicating its effectiveness in detecting objects from a bird’s eye view perspective.
MonoDETRNext-A also demonstrates competitive performance, particularly in the validation set where it achieves the highest $AP_{3D}$ and $AP_{BEV}$ scores among all models.
Furthermore, both MonoDETRNext-F and MonoDETRNext-A exhibit respectable inference speeds, with MonoDETRNext-F achieving the highest $FPS_{bz=1}$ among all models.
4.3 Ablation Studies
We have validated the effectiveness of each component and strategy, and reported the $AP_{3D}$ for the car category on the KITTI validation benchmark.
Our study evaluates different strategies for object query generation, as depicted in Table 2(a). "L-center assign learnable embeddings," "enc-box," and "enc-center" refer to the anchors derived from backbone features, generated separately by two heads and contributing to individual loss calculation. Additionally, "enc-output" signifies six-dimensional anchors produced by a single head, also participating in loss calculation. Among these strategies, the proposed effective depth-aware query generation demonstrates notable performance, as evidenced by superior results across Easy, Moderate, and Hard stages compared to alternative methods.
Table 2(b) illustrates the performance of MonoDETRNext-F with different vision encoders, including the "hybrid" encoder featured in Figure 2. Notably, the "hybrid" encoder demonstrates superior performance across Easy, Moderate, and Hard categories compared to other configurations. Interestingly, even with fewer layers, the "hybrid-L" encoder maintains competitive performance, suggesting that a lower layer count may be more effective. This observation prompts speculation: the efficacy of fewer layers could stem from the relatively limited KITTI dataset or indicate that a single layer may constitute a more rational design choice. Further analysis is warranted to discern whether this trend holds across diverse datasets and to explore potential implications for architectural design in 3d object detection tasks.
In Table 2(d), the impact of different depth positional encodings on MonoDETRNext-F is examined. "3D sin/cos" and "2D sin/cos" utilize sinusoidal functions for encoding, with the former incorporating all six output dimensions. Notably, "3D sin/cos" outperforms other settings across Easy, Moderate, and Hard categories. This suggests that the precise encoding of depth information through sinusoidal functions significantly benefits object detection. Conversely, "Meter-wise" and "k-bin" positional encodings yield comparatively lower performance, indicating less effective depth representation in these configurations. Further investigation into the mechanisms behind these encoding methods is warranted to optimize depth-aware object detection algorithms.
In Table 2(c), $N_{1}$, $N_{2}$, and $N_{3}$ signify the number of SDC modules, as depicted in Figure 2 .
(a) Performance of MonoDETRNext-F at different 3D object query generation strategy
Stage Easy Mod. Hard
L-center+enc-box 30.28 24.77 20.38
enc-center+enc-box 23.89 19.49 17.97
enc-output 23.80 18.96 16.47
w/o 22.97 17.01 15.45
(b) Performance of MonoDETRNext-F under different vision encoder
Model Easy Mod. Hard
hybrid 30.28 24.77 20.38
hybrid-L 29.42 22.01 18.99
MonoDETR’s 28.20 22.06 18.32
RT-DETR’s 26.42 20.45 16.78
(c) Effect of Different Scales of Accurate Depth Predictors on MonoDETRNext-A Performance
$N_{1},N_{2},N_{3}$ Easy Moderate Hard
2, 2, 4 33.44 25.75 22.40
1, 1, 2 30.26 24.99 22.58
3, 3 ,6 32.45 23.85 21.85
4, 4, 6 31.78 23.56 20.78
(d) The impact of different depth positional encodings of MonoDETRNext-F.
Settings Easy Mod. Hard
3D sin/cos 30.28 24.77 20.38
2D sin/cos 28.21 21.52 18.77
Meter-wise 26.29 20.04 18.13
k-bin 25.20 18.51 16.85
Table 2: Ablation Studies on Our Proposed MonoDETRNext on the KITTI validation benchmark. The specific meanings of the table can be referred to in Section 4.3.
5 Conclusion and Limitations
Conclusion: This paper introduces a novel monocular vision-based approach for 3D object detection. Leveraging advancements from the 2D detection domain, we propose the efficient and precise MonoDETRNext. Building upon the groundwork laid by MonoDETR, we introduce two variants: MonoDETRNext-F prioritizing speed, and MonoDETRNext-A emphasizing accuracy. Our methodology encompasses the development of an efficient hybrid vision encoder, enhancements to depth prediction mechanisms, and refinements in object query generation. Through comprehensive performance evaluation, we establish the superiority of our model over existing methods. By optimizing both accuracy and computational efficiency, MonoDETRNext sets a new benchmark in monocular 3D object detection, facilitating future research and applications across diverse real-world scenarios.
Limitations: Despite the substantial advancements achieved by MonoDETRNext in enhancing the accuracy and computational efficiency of monocular 3D object detection, certain limitations persist. Due to the inherent constraints of monocular vision methods, there remains a notable discrepancy in accuracy and performance compared to approaches employing multi-view methodologies or sensor fusion techniques, such as the integration of LiDAR with cameras.